# PICore: Physics-Informed Unsupervised Coreset Selection for Data Efficient Neural Operator Training

### v1

@@ 6.1 Main Results
While PICore offers strong performance and efficiency, one tradeoff is that the absolute test accuracy relative to training on $100\%$ of the data is lower. For example, on the Advection dataset with FNO, the 100% training baseline yields an NRMSE of $2.13\times 10^{-2}$, while PICore at 20% yields $3.77\times 10^{-2}$. However, this is an inherent tradeoff for all coreset selection algorithms, as the selected coreset simply contains less information for training. Additionally, this is not specific to PICore, as similar reductions in accuracy hold for supervised coreset selection. In practice, one may want to select a higher selection percentage, such as $40\%$, which would yield higher accuracy ($2.86\times 10^{-2}$) while still maintaining a competitive efficiency gain (2.55×2.$5$).
Operator Method Algorithm 20.0% 30.0% 40.0% 60.0% 80.0% 100.0%
FNO Supervised craig $3.45\times 10^{-2}$ $3.12\times 10^{-2}$ $2.73\times 10^{-2}$ $2.49\times 10^{-2}$ $2.30\times 10^{-2}$ $2.13\times 10^{-2}$
gradmatch $3.48\times 10^{-2}$ $2.96\times 10^{-2}$ $2.64\times 10^{-2}$ $2.43\times 10^{-2}$ $2.22\times 10^{-2}$ $2.13\times 10^{-2}$
adacore $4.76\times 10^{-2}$ $3.94\times 10^{-2}$ $3.21\times 10^{-2}$ $2.61\times 10^{-2}$ $2.34\times 10^{-2}$ $2.13\times 10^{-2}$
el2n $3.35\times 10^{-2}$ $2.92\times 10^{-2}$ $2.73\times 10^{-2}$ $2.47\times 10^{-2}$ $2.33\times 10^{-2}$ $2.13\times 10^{-2}$
graNd $3.42\times 10^{-2}$ $2.96\times 10^{-2}$ $2.64\times 10^{-2}$ $2.42\times 10^{-2}$ $2.25\times 10^{-2}$ $2.13\times 10^{-2}$
Acceleration 4.76×4.76$\times$4.76 × 3.19×3.19$\times$3.19 × 2.49×2.49$\times$2.49 × 1.69×1.69$\times$1.69 × 1.27×1.27$\times$1.27 × 1.00×1.$0$
PICore craig $3.66\times 10^{-2}$ $3.28\times 10^{-2}$ $2.94\times 10^{-2}$ $2.56\times 10^{-2}$ $2.32\times 10^{-2}$ $2.13\times 10^{-2}$
gradmatch $3.38\times 10^{-2}$ $3.22\times 10^{-2}$ $2.70\times 10^{-2}$ $2.44\times 10^{-2}$ $2.20\times 10^{-2}$ $2.13\times 10^{-2}$
adacore $5.08\times 10^{-2}$ $4.00\times 10^{-2}$ $3.26\times 10^{-2}$ $2.70\times 10^{-2}$ $2.40\times 10^{-2}$ $2.13\times 10^{-2}$
el2n $3.29\times 10^{-2}$ $3.03\times 10^{-2}$ $2.71\times 10^{-2}$ $2.49\times 10^{-2}$ $2.35\times 10^{-2}$ $2.13\times 10^{-2}$
graNd $3.46\times 10^{-2}$ $3.04\times 10^{-2}$ $2.69\times 10^{-2}$ $2.40\times 10^{-2}$ $2.25\times 10^{-2}$ $2.13\times 10^{-2}$
Acceleration 5.10×5.10$\times$5.10 × 3.30×3.30$\times$3.30 × 2.55×2.$5$ 1.70×1.70$\times$1.70 × 1.27×1.27$\times$1.27 × 1.00×1.$0$
UNO Supervised craig $1.57\times 10^{-1}$ $1.50\times 10^{-1}$ $1.43\times 10^{-1}$ $1.27\times 10^{-1}$ $8.45\times 10^{-2}$ $7.24\times 10^{-2}$
gradmatch $1.55\times 10^{-1}$ $1.48\times 10^{-1}$ $1.42\times 10^{-1}$ $1.17\times 10^{-1}$ $8.69\times 10^{-2}$ $7.24\times 10^{-2}$
adacore $2.21\times 10^{-1}$ $1.77\times 10^{-1}$ $1.59\times 10^{-1}$ $1.42\times 10^{-1}$ $1.00\times 10^{-1}$ $7.24\times 10^{-2}$
el2n $1.54\times 10^{-1}$ $1.47\times 10^{-1}$ $1.42\times 10^{-1}$ $1.33\times 10^{-1}$ $9.68\times 10^{-2}$ $7.24\times 10^{-2}$
graNd $1.53\times 10^{-1}$ $1.47\times 10^{-1}$ $1.41\times 10^{-1}$ $1.27\times 10^{-1}$ $9.45\times 10^{-2}$ $7.24\times 10^{-2}$
Acceleration 4.91×4.91$\times$4.91 × 3.28×3.28$\times$3.28 × 2.51×2.51$\times$2.51 × 1.70×1.70$\times$1.70 × 1.27×1.27$\times$1.27 × 1.00×1.$0$
PICore craig $1.60\times 10^{-1}$ $1.48\times 10^{-1}$ $1.44\times 10^{-1}$ $1.29\times 10^{-1}$ $9.67\times 10^{-2}$ $7.24\times 10^{-2}$
gradmatch $1.55\times 10^{-1}$ $1.47\times 10^{-1}$ $1.43\times 10^{-1}$ $1.26\times 10^{-1}$ $9.06\times 10^{-2}$ $7.24\times 10^{-2}$
adacore $2.25\times 10^{-1}$ $1.71\times 10^{-1}$ $1.47\times 10^{-1}$ $1.31\times 10^{-1}$ $1.24\times 10^{-1}$ $7.24\times 10^{-2}$
el2n $1.52\times 10^{-1}$ $1.45\times 10^{-1}$ $1.42\times 10^{-1}$ $1.30\times 10^{-1}$ $9.89\times 10^{-2}$ $7.24\times 10^{-2}$
graNd $1.54\times 10^{-1}$ $1.46\times 10^{-1}$ $1.42\times 10^{-1}$ $1.33\times 10^{-1}$ $9.04\times 10^{-2}$ $7.24\times 10^{-2}$
Acceleration 5.07×5.07$\times$5.07 × $3\times3$ 2.53×2.53$\times$2.53 × 1.69×1.69$\times$1.69 × 1.26×1.26$\times$1.26 × 1.00×1.$0$
Table 1: Test NRMSE and relative training acceleration on the Advection dataset at resolution 64 across varying coreset percentages (20%–100%) between supervised and PICore-based coreset selection methods using both FNO and UNO architectures.
Operator Method Algorithm 20.0% 30.0% 40.0% 60.0% 80.0% 100.0%
FNO Supervised craig $1.42\times 10^{-2}$ $1.44\times 10^{-2}$ $8.75\times 10^{-3}$ $5.54\times 10^{-3}$ $4.26\times 10^{-3}$ $3.57\times 10^{-3}$
gradmatch $1.71\times 10^{-2}$ $1.12\times 10^{-2}$ $7.68\times 10^{-3}$ $5.24\times 10^{-3}$ $4.13\times 10^{-3}$ $3.57\times 10^{-3}$
adacore $2.65\times 10^{-2}$ $2.46\times 10^{-2}$ $1.61\times 10^{-2}$ $8.27\times 10^{-3}$ $4.74\times 10^{-3}$ $3.57\times 10^{-3}$
el2n $1.67\times 10^{-2}$ $1.16\times 10^{-2}$ $8.25\times 10^{-3}$ $5.61\times 10^{-3}$ $4.36\times 10^{-3}$ $3.57\times 10^{-3}$
graNd $1.67\times 10^{-2}$ $1.27\times 10^{-2}$ $8.50\times 10^{-3}$ $5.26\times 10^{-3}$ $4.24\times 10^{-3}$ $3.57\times 10^{-3}$
Acceleration 3.26×3.26$\times$3.26 × 2.50×2.50$\times$2.50 × 2.09×2.09$\times$2.09 × 1.54×1.54$\times$1.54 × 1.22×1.$2$ 1.00×1.$0$
PICore craig $1.93\times 10^{-2}$ $1.56\times 10^{-2}$ $1.10\times 10^{-2}$ $6.68\times 10^{-3}$ $4.40\times 10^{-3}$ $3.57\times 10^{-3}$
gradmatch $1.91\times 10^{-2}$ $1.15\times 10^{-2}$ $8.08\times 10^{-3}$ $5.61\times 10^{-3}$ $4.24\times 10^{-3}$ $3.57\times 10^{-3}$
adacore $2.86\times 10^{-2}$ $2.32\times 10^{-2}$ $1.46\times 10^{-2}$ $8.07\times 10^{-3}$ $4.90\times 10^{-3}$ $3.57\times 10^{-3}$
el2n $1.81\times 10^{-2}$ $1.12\times 10^{-2}$ $8.07\times 10^{-3}$ $5.49\times 10^{-3}$ $4.07\times 10^{-3}$ $3.57\times 10^{-3}$
graNd $1.87\times 10^{-2}$ $1.26\times 10^{-2}$ $8.15\times 10^{-3}$ $5.50\times 10^{-3}$ $4.08\times 10^{-3}$ $3.57\times 10^{-3}$
Acceleration 5.05×5.05$\times$5.05 × 3.30×3.30$\times$3.30 × 2.53×2.53$\times$2.53 × 1.69×1.69$\times$1.69 × 1.26×1.26$\times$1.26 × 1.00×1.$0$
UNO Supervised craig $3.01\times 10^{-2}$ $2.49\times 10^{-2}$ $2.20\times 10^{-2}$ $1.78\times 10^{-2}$ $1.54\times 10^{-2}$ $1.70\times 10^{-2}$
gradmatch $2.93\times 10^{-2}$ $2.42\times 10^{-2}$ $2.08\times 10^{-2}$ $1.73\times 10^{-2}$ $1.54\times 10^{-2}$ $1.70\times 10^{-2}$
adacore $4.67\times 10^{-2}$ $3.61\times 10^{-2}$ $3.05\times 10^{-2}$ $2.13\times 10^{-2}$ $1.69\times 10^{-2}$ $1.70\times 10^{-2}$
el2n $3.00\times 10^{-2}$ $2.55\times 10^{-2}$ $2.15\times 10^{-2}$ $1.79\times 10^{-2}$ $1.56\times 10^{-2}$ $1.70\times 10^{-2}$
graNd $2.99\times 10^{-2}$ $2.59\times 10^{-2}$ $2.19\times 10^{-2}$ $1.80\times 10^{-2}$ $1.55\times 10^{-2}$ $1.70\times 10^{-2}$
Acceleration 3.76×3.76$\times$3.76 × 2.76×2.76$\times$2.76 × $2\times2$ 1.58×1.58$\times$1.58 × 1.22×1.$2$ 1.00×1.$0$
PICore craig $3.28\times 10^{-2}$ $3.01\times 10^{-2}$ $2.33\times 10^{-2}$ $1.90\times 10^{-2}$ $1.57\times 10^{-2}$ $1.70\times 10^{-2}$
gradmatch $2.96\times 10^{-2}$ $2.43\times 10^{-2}$ $2.08\times 10^{-2}$ $1.75\times 10^{-2}$ $1.56\times 10^{-2}$ $1.70\times 10^{-2}$
adacore $5.09\times 10^{-2}$ $3.96\times 10^{-2}$ $3.17\times 10^{-2}$ $2.21\times 10^{-2}$ $1.72\times 10^{-2}$ $1.70\times 10^{-2}$
el2n $2.95\times 10^{-2}$ $2.47\times 10^{-2}$ $2.10\times 10^{-2}$ $1.73\times 10^{-2}$ $1.51\times 10^{-2}$ $1.70\times 10^{-2}$
graNd $2.84\times 10^{-2}$ $2.36\times 10^{-2}$ $2.06\times 10^{-2}$ $1.72\times 10^{-2}$ $1.57\times 10^{-2}$ $1.70\times 10^{-2}$
Acceleration 5.02×5.02$\times$5.02 × 3.31×3.31$\times$3.31 × 2.51×2.51$\times$2.51 × 1.68×1.68$\times$1.68 × 1.25×1.25$\times$1.25 × 1.00×1.$0$
Table 2: Test NRMSE and relative training acceleration on the Burgers dataset at resolution 64 across varying coreset percentages (20%–100%) between supervised and PICore-based coreset selection methods using both FNO and UNO architectures.
Operator Method Algorithm 20.0% 30.0% 40.0% 60.0% 80.0% 100.0%
FNO Supervised craig $1.26\times 10^{-1}$ $1.15\times 10^{-1}$ $1.01\times 10^{-1}$ $7.87\times 10^{-2}$ $6.67\times 10^{-2}$ $5.96\times 10^{-2}$
gradmatch $1.25\times 10^{-1}$ $1.09\times 10^{-1}$ $9.48\times 10^{-2}$ $7.62\times 10^{-2}$ $6.71\times 10^{-2}$ $5.96\times 10^{-2}$
adacore $1.62\times 10^{-1}$ $1.34\times 10^{-1}$ $1.12\times 10^{-1}$ $8.64\times 10^{-2}$ $7.01\times 10^{-2}$ $5.96\times 10^{-2}$
el2n $1.26\times 10^{-1}$ $1.07\times 10^{-1}$ $9.43\times 10^{-2}$ $7.83\times 10^{-2}$ $6.59\times 10^{-2}$ $5.96\times 10^{-2}$
graNd $1.27\times 10^{-1}$ $1.09\times 10^{-1}$ $9.32\times 10^{-2}$ $7.71\times 10^{-2}$ $6.61\times 10^{-2}$ $5.96\times 10^{-2}$
Acceleration 1.98×1.98$\times$1.98 × 1.76×1.76$\times$1.76 × 1.59×1.59$\times$1.59 × 1.33×1.$3$ 1.14×1.14$\times$1.14 × 1.00×1.$0$
PICore craig $1.36\times 10^{-1}$ $1.20\times 10^{-1}$ $1.05\times 10^{-1}$ $8.20\times 10^{-2}$ $6.85\times 10^{-2}$ $5.96\times 10^{-2}$
gradmatch $1.33\times 10^{-1}$ $1.13\times 10^{-1}$ $9.97\times 10^{-2}$ $7.92\times 10^{-2}$ $6.68\times 10^{-2}$ $5.96\times 10^{-2}$
adacore $1.66\times 10^{-1}$ $1.33\times 10^{-1}$ $1.11\times 10^{-1}$ $8.00\times 10^{-2}$ $6.88\times 10^{-2}$ $5.96\times 10^{-2}$
el2n $1.25\times 10^{-1}$ $1.12\times 10^{-1}$ $9.44\times 10^{-2}$ $7.77\times 10^{-2}$ $6.84\times 10^{-2}$ $5.96\times 10^{-2}$
graNd $1.32\times 10^{-1}$ $1.11\times 10^{-1}$ $9.75\times 10^{-2}$ $7.82\times 10^{-2}$ $6.81\times 10^{-2}$ $5.96\times 10^{-2}$
Acceleration 5.01×5.01$\times$5.01 × 3.32×3.32$\times$3.32 × 2.50×2.50$\times$2.50 × 1.67×1.67$\times$1.67 × 1.25×1.25$\times$1.25 × 1.00×1.$0$
UNO Supervised craig $1.35\times 10^{-1}$ $1.17\times 10^{-1}$ $1.05\times 10^{-1}$ $9.10\times 10^{-2}$ $7.91\times 10^{-2}$ $6.68\times 10^{-2}$
gradmatch $1.28\times 10^{-1}$ $1.14\times 10^{-1}$ $9.84\times 10^{-2}$ $8.60\times 10^{-2}$ $7.70\times 10^{-2}$ $6.68\times 10^{-2}$
adacore $1.80\times 10^{-1}$ $1.50\times 10^{-1}$ $1.26\times 10^{-1}$ $9.73\times 10^{-2}$ $8.47\times 10^{-2}$ $6.68\times 10^{-2}$
el2n $1.39\times 10^{-1}$ $1.15\times 10^{-1}$ $1.01\times 10^{-1}$ $8.87\times 10^{-2}$ $7.87\times 10^{-2}$ $6.68\times 10^{-2}$
graNd $1.31\times 10^{-1}$ $1.15\times 10^{-1}$ $1.01\times 10^{-1}$ $8.70\times 10^{-2}$ $7.87\times 10^{-2}$ $6.68\times 10^{-2}$
Acceleration 2.24×2.24$\times$2.24 × 1.94×1.94$\times$1.94 × 1.71×1.71$\times$1.71 × 1.39×1.39$\times$1.39 × 1.17×1.17$\times$1.17 × 1.00×1.$0$
PICore craig $1.49\times 10^{-1}$ $1.32\times 10^{-1}$ $1.14\times 10^{-1}$ $9.31\times 10^{-2}$ $8.00\times 10^{-2}$ $6.68\times 10^{-2}$
gradmatch $1.29\times 10^{-1}$ $1.10\times 10^{-1}$ $9.87\times 10^{-2}$ $8.44\times 10^{-2}$ $7.69\times 10^{-2}$ $6.68\times 10^{-2}$
adacore $1.84\times 10^{-1}$ $1.51\times 10^{-1}$ $1.25\times 10^{-1}$ $9.86\times 10^{-2}$ $8.21\times 10^{-2}$ $6.68\times 10^{-2}$
el2n $1.33\times 10^{-1}$ $1.14\times 10^{-1}$ $9.83\times 10^{-2}$ $8.45\times 10^{-2}$ $7.55\times 10^{-2}$ $6.68\times 10^{-2}$
graNd $1.28\times 10^{-1}$ $1.12\times 10^{-1}$ $9.67\times 10^{-2}$ $8.42\times 10^{-2}$ $7.61\times 10^{-2}$ $6.68\times 10^{-2}$
Acceleration 5.01×5.01$\times$5.01 × 3.32×3.32$\times$3.32 × 2.50×2.50$\times$2.50 × 1.67×1.67$\times$1.67 × 1.25×1.25$\times$1.25 × 1.00×1.$0$
Table 3: Test NRMSE and relative training acceleration on the Darcy dataset at resolution 64 across varying coreset percentages (20%–100%) between supervised and PICore-based coreset selection methods using both FNO and UNO architectures.
[…]

### v2

@@ 6.1 Main Results
While PICore offers strong performance and efficiency, one tradeoff is that the absolute test accuracy relative to training on $100\%$ of the data is lower. For example, on the Advection dataset with FNO, the 100% training baseline yields an NRMSE of $2.13\times 10^{-2}$, while PICore at 20% yields $3.77\times 10^{-2}$. However, this is an inherent tradeoff for all coreset selection algorithms, as the selected coreset simply contains less information for training. Additionally, this is not specific to PICore, as similar reductions in accuracy hold for supervised coreset selection. In practice, one may want to select a higher selection percentage, such as $40\%$, which would yield higher accuracy ($2.69\times 10^{-2}$) while still maintaining a competitive gain ($2.54\times$).
Table 1: Advection NRMSE at resolution 64
Operator Method 20.0% 30.0% 40.0% 60.0% 80.0% 100.0%
FNO Physics-Informed $8.43\pm 0.03\times 10^{-1}$ $8.57\pm 0.03\times 10^{-1}$ $8.73\pm 0.04\times 10^{-1}$ $8.98\pm 0.02\times 10^{-1}$ $9.14\pm 0.02\times 10^{-1}$ $9.26\pm 0.03\times 10^{-1}$
$(4.72\times)$ $(3.06\times)$ $(2.38\times)$ $(1.59\times)$ $(1.19\times)$ $(0.96\times)$
Random $\textbf{3.39}\pm 0.07\times 10^{-2}$ $\textbf{2.89}\pm 0.03\times 10^{-2}$ $\underline{2.68}\pm 0.02\times 10^{-2}$ $2.47\pm 0.03\times 10^{-2}$ $37\pm 0.04\times 10^{-2} 2.2$$22\pm 0.05\times 10^{-2}$
$(5.10\times)$ $(3.32\times)$ $(2.56\times)$ $(1.72\times)$ $(1.28\times)$ $(1.00\times)$
Active Learning $8.32\pm 0.58\times 10^{-2}$ $6.29\pm 0.40\times 10^{-2}$ $4.78\pm 0.27\times 10^{-2}$ $3.51\pm 0.16\times 10^{-2}$ $96\pm 0.07\times 10^{-2} 2.2$$22\pm 0.05\times 10^{-2}$
$(5.04\times)$ $(3.28\times)$ $(2.52\times)$ $(1.69\times)$ $(1.26\times)$ $(1.00\times)$
Supervised (graNd) $\underline{3.42}\pm 0.12\times 10^{-2}$ $\underline{2.96}\pm 0.09\times 10^{-2}$ $\textbf{2.64}\pm 0.03\times 10^{-2}$ $\underline{2.42}\pm 0.03\times 10^{-2}$ $\underline{2.25}\pm 0.02\times 10^{-2}$ $2.22\pm 0.05\times 10^{-2}$
$(4.70\times)$ $(3.15\times)$ $(2.45\times)$ $(1.66\times)$ $(1.26\times)$ $(1.00\times)$
PICore (graNd) $3.46\pm 0.13\times 10^{-2}$ $3.04\pm 0.15\times 10^{-2}$ $2.69\pm 0.05\times 10^{-2}$ $\textbf{2.40}\pm 0.04\times 10^{-2}$ $\textbf{2.25}\pm 0.04\times 10^{-2}$ $2.22\pm 0.05\times 10^{-2}$
$(5.06\times)$ $(3.27\times)$ $(2.54\times)$ $(1.68\times)$ $(1.26\times)$ $(1.00\times)$
UNO Physics-Informed $8.07\pm 0.05\times 10^{-1}$ $8.19\pm 0.06\times 10^{-1}$ $8.28\pm 0.05\times 10^{-1}$ $8.46\pm 0.18\times 10^{-1}$ $9.09\pm 0.37\times 10^{-1}$ $9.25\pm 0.30\times 10^{-1}$
$(4.87\times)$ $(3.21\times)$ $(2.46\times)$ $(1.65\times)$ $(1.23\times)$ $(0.98\times)$
Random $1.59\pm 0.02\times 10^{-1}$ $1.50\pm 0.01\times 10^{-1}$ $1.44\pm 0.007\times 10^{-1}$ $1.42\pm 0.12\times 10^{-1}$ $1.32\pm 0.12\times 10^{-1}$ $7.27\pm 0.28\times 10^{-2}$
$(5.08\times)$ $(3.35\times)$ $(2.55\times)$ $(1.70\times)$ $(1.28\times)$ $(1.00\times)$
Active Learning $1.96\pm 0.04\times 10^{-1}$ $1.59\pm 0.05\times 10^{-1}$ $\textbf{9.20}\pm 0.73\times 10^{-2}$ $\textbf{7.49}\pm 0.47\times 10^{-2}$ $\textbf{6.83}\pm 0.05\times 10^{-2}$ $7.27\pm 0.28\times 10^{-2}$
$(5.05\times)$ $(3.32\times)$ $(2.52\times)$ $(1.68\times)$ $(1.26\times)$ $(1.00\times)$
Supervised (gradmatch) $\textbf{1.55}\pm 0.02\times 10^{-1}$ $\underline{1.48}\pm 0.01\times 10^{-1}$ $\underline{1.42}\pm 0.02\times 10^{-1}$ $\underline{1.17}\pm 0.14\times 10^{-1}$ $\underline{8.69}\pm 1.23\times 10^{-2}$ $7.27\pm 0.28\times 10^{-2}$
$(4.84\times)$ $(3.23\times)$ $(2.47\times)$ $(1.67\times)$ $(1.25\times)$ $(1.00\times)$
PICore (gradmatch) $\textbf{1.55}\pm 0.01\times 10^{-1}$ $\textbf{1.47}\pm 0.01\times 10^{-1}$ $1.43\pm 0.008\times 10^{-1}$ $1.26\pm 0.09\times 10^{-1}$ $9.06\pm 1.07\times 10^{-2}$ $7.27\pm 0.28\times 10^{-2}$
$(5.07\times)$ $(3.34\times)$ $(2.53\times)$ $(1.69\times)$ $(1.26\times)$ $(1.00\times)$
Table 2: Burgers NRMSE at resolution 64
Operator Method 20.0% 30.0% 40.0% 60.0% 80.0% 100.0%
FNO Physics-Informed $5.26\pm 0.12\times 10^{-1}$ $4.55\pm 0.06\times 10^{-1}$ $4.36\pm 0.07\times 10^{-1}$ $4.23\pm 0.03\times 10^{-1}$ $4.19\pm 0.03\times 10^{-1}$ $4.12\pm 0.01\times 10^{-1}$
$(5.36\times)$ $(3.45\times)$ $(2.70\times)$ $(1.80\times)$ $(1.34\times)$ $(1.07\times)$
Random $85\pm 0.09\times 10^{-2} 1.1$$18\pm 0.06\times 10^{-2}$ $8.23\pm 0.21\times 10^{-3}$ $5.82\pm 0.21\times 10^{-3}$ $4.75\pm 0.13\times 10^{-3}$ $3.95\pm 0.10\times 10^{-3}$
$(5.07\times)$ $(3.31\times)$ $(2.56\times)$ $(1.72\times)$ $(1.28\times)$ $(1.00\times)$
Active Learning $8.76\pm 2.52\times 10^{-2}$ $4.57\pm 0.95\times 10^{-2}$ $3.30\pm 0.57\times 10^{-2}$ $2.06\pm 0.07\times 10^{-2}$ $1.37\pm 0.18\times 10^{-2}$ $3.95\pm 0.10\times 10^{-3}$
$(5.03\times)$ $(3.25\times)$ $(2.52\times)$ $(1.68\times)$ $(1.26\times)$ $(1.00\times)$
Supervised (gradmatch) $\textbf{1.71}\pm 0.16\times 10^{-2}$ $\underline{1.12}\pm 0.09\times 10^{-2}$ $\textbf{7.68}\pm 0.29\times 10^{-3}$ $\textbf{5.24}\pm 0.14\times 10^{-3}$ $\underline{4.13}\pm 0.08\times 10^{-3}$ $3.95\pm 0.10\times 10^{-3}$
$(3.28\times)$ $(2.52\times)$ $(2.11\times)$ $(1.55\times)$ $(1.22\times)$ $(1.00\times)$
PICore (el2n) $\underline{1.81}\pm 0.08\times 10^{-2}$ $\textbf{1.12}\pm 0.07\times 10^{-2}$ $\underline{8.07}\pm 0.33\times 10^{-3}$ $\underline{5.49}\pm 0.08\times 10^{-3}$ $\textbf{4.07}\pm 0.10\times 10^{-3}$ $3.95\pm 0.10\times 10^{-3}$
$(5.05\times)$ $(3.30\times)$ $(2.53\times)$ $(1.68\times)$ $(1.26\times)$ $(1.00\times)$
UNO Physics-Informed $4.82\pm 0.07\times 10^{-1}$ $56\pm 0.06\times 10^{-1} 4.4$$48\pm 0.06\times 10^{-1}$ $4.61\pm 0.03\times 10^{-1}$ $4.63\pm 0.02\times 10^{-1}$ $4.65\pm 0.04\times 10^{-1}$
$(5.28\times)$ $(3.46\times)$ $(2.66\times)$ $(1.77\times)$ $(1.33\times)$ $(1.06\times)$
Random $\underline{2.92}\pm 0.05\times 10^{-2}$ $55\pm 0.05\times 10^{-2} 2.2$$25\pm 0.05\times 10^{-2}$ $1.83\pm 0.03\times 10^{-2}$ $1.58\pm 0.01\times 10^{-2}$ $1.49\pm 0.04\times 10^{-2}$
$(4.99\times)$ $(3.34\times)$ $(2.55\times)$ $(1.70\times)$ $(1.27\times)$ $(1.00\times)$
Active Learning $5.42\pm 0.41\times 10^{-2}$ $4.12\pm 0.11\times 10^{-2}$ $3.62\pm 0.20\times 10^{-2}$ $3.03\pm 0.23\times 10^{-2}$ $2.51\pm 0.13\times 10^{-2}$ $1.49\pm 0.04\times 10^{-2}$
$(5.01\times)$ $(3.30\times)$ $(2.51\times)$ $(1.68\times)$ $(1.26\times)$ $(1.00\times)$
Supervised (gradmatch) $2.93\pm 0.10\times 10^{-2}$ $\underline{2.42}\pm 0.06\times 10^{-2}$ $\underline{2.08}\pm 0.05\times 10^{-2}$ $\underline{1.73}\pm 0.03\times 10^{-2}$ $\textbf{1.54}\pm 0.02\times 10^{-2}$ $1.49\pm 0.04\times 10^{-2}$
$(3.77\times)$ $(2.77\times)$ $(2.23\times)$ $(1.59\times)$ $(1.23\times)$ $(1.00\times)$
PICore (graNd) $\textbf{2.84}\pm 0.05\times 10^{-2}$ $\textbf{2.36}\pm 0.05\times 10^{-2}$ $\textbf{2.06}\pm 0.04\times 10^{-2}$ $\textbf{1.72}\pm 0.05\times 10^{-2}$ $\underline{1.57}\pm 0.03\times 10^{-2}$ $1.49\pm 0.04\times 10^{-2}$
$(5.05\times)$ $(3.33\times)$ $(2.52\times)$ $(1.69\times)$ $(1.26\times)$ $(1.00\times)$
Table 3: Darcy NRMSE at resolution 64
Operator Method 20.0% 30.0% 40.0% 60.0% 80.0% 100.0%
FNO Physics-Informed $1.46\pm 0.003\times 10^{0}$ $1.47\pm 0.003\times 10^{0}$ $1.47\pm 0.002\times 10^{0}$ $1.47\pm 0.001\times 10^{0}$ $1.47\pm 0.002\times 10^{0}$ $1.48\pm 0.001\times 10^{0}$
$(7.43\times)$ $(4.96\times)$ $(3.75\times)$ $(2.50\times)$ $(1.88\times)$ $(1.51\times)$
Random $34\pm 0.03\times 10^{-1} 1.1$$15\pm 0.01\times 10^{-1}$ $9.99\pm 0.11\times 10^{-2}$ $7.94\pm 0.04\times 10^{-2}$ $7.07\pm 0.16\times 10^{-2}$ $6.18\pm 0.09\times 10^{-2}$
$(5.00\times)$ $(3.36\times)$ $(2.53\times)$ $(1.69\times)$ $(1.27\times)$ $(1.00\times)$
Active Learning $2.01\pm 0.16\times 10^{-1}$ $1.58\pm 0.08\times 10^{-1}$ $1.25\pm 0.06\times 10^{-1}$ $8.94\pm 0.33\times 10^{-2}$ $7.19\pm 0.25\times 10^{-2}$ $6.18\pm 0.09\times 10^{-2}$
$(4.99\times)$ $(3.31\times)$ $(2.50\times)$ $(1.66\times)$ $(1.25\times)$ $(1.00\times)$
Supervised (el2n) $\underline{1.26}\pm 0.01\times 10^{-1}$ $\textbf{1.07}\pm 0.007\times 10^{-1}$ $\textbf{9.43}\pm 0.09\times 10^{-2}$ $\underline{7.83}\pm 0.18\times 10^{-2}$ $\textbf{6.59}\pm 0.09\times 10^{-2}$ $6.18\pm 0.09\times 10^{-2}$
$(1.98\times)$ $(1.76\times)$ $(1.59\times)$ $(1.33\times)$ $(1.14\times)$ $(1.00\times)$
PICore (el2n) $\textbf{1.25}\pm 0.02\times 10^{-1}$ $\underline{1.12}\pm 0.02\times 10^{-1}$ $\underline{9.44}\pm 0.12\times 10^{-2}$ $\textbf{7.77}\pm 0.18\times 10^{-2}$ $\underline{6.84}\pm 0.18\times 10^{-2}$ $6.18\pm 0.09\times 10^{-2}$
$(5.00\times)$ $(3.32\times)$ $(2.50\times)$ $(1.67\times)$ $(1.25\times)$ $(1.00\times)$
UNO Physics-Informed $1.42\pm 0.002\times 10^{0}$ $1.42\pm 0.003\times 10^{0}$ $1.42\pm 0.003\times 10^{0}$ $1.43\pm 0.001\times 10^{0}$ $1.42\pm 0.002\times 10^{0}$ $1.43\pm 0.003\times 10^{0}$
$(7.05\times)$ $(4.71\times)$ $(3.55\times)$ $(2.37\times)$ $(1.78\times)$ $(1.46\times)$
Random $1.45\pm 0.02\times 10^{-1}$ $22\pm 0.03\times 10^{-1} 1.1$$10\pm 0.02\times 10^{-1}$ $9.23\pm 0.22\times 10^{-2}$ $8.78\pm 0.30\times 10^{-2}$ $7.57\pm 0.13\times 10^{-2}$
$(5.03\times)$ $(3.37\times)$ $(2.53\times)$ $(1.69\times)$ $(1.27\times)$ $(1.00\times)$
Active Learning $1.87\pm 0.14\times 10^{-1}$ $1.54\pm 0.08\times 10^{-1}$ $1.27\pm 0.06\times 10^{-1}$ $1.02\pm 0.03\times 10^{-1}$ $8.63\pm 0.19\times 10^{-2}$ $7.57\pm 0.13\times 10^{-2}$
$(5.04\times)$ $(3.35\times)$ $(2.52\times)$ $(1.68\times)$ $(1.26\times)$ $(1.00\times)$
Supervised (gradmatch) $\textbf{1.28}\pm 0.03\times 10^{-1}$ $\underline{1.14}\pm 0.01\times 10^{-1}$ $\underline{9.84}\pm 0.16\times 10^{-2}$ $\underline{8.60}\pm 0.11\times 10^{-2}$ $\underline{7.70}\pm 0.10\times 10^{-2}$ $7.57\pm 0.13\times 10^{-2}$
$(2.23\times)$ $(1.93\times)$ $(1.71\times)$ $(1.38\times)$ $(1.16\times)$ $(1.00\times)$
PICore (graNd) $\textbf{1.28}\pm 0.03\times 10^{-1}$ $\textbf{1.12}\pm 0.01\times 10^{-1}$ $\textbf{9.67}\pm 0.16\times 10^{-2}$ $\textbf{8.42}\pm 0.14\times 10^{-2}$ $\textbf{7.61}\pm 0.11\times 10^{-2}$ $7.57\pm 0.13\times 10^{-2}$
$(5.01\times)$ $(3.33\times)$ $(2.50\times)$ $(1.67\times)$ $(1.25\times)$ $(1.00\times)$
Table 4: Navier Stokes Incompressible NRMSE at resolution 64
Operator Method 20.0% 30.0% 40.0% 60.0% 80.0% 100.0%
FNO Physics-Informed $1\pm 0.001\times 10^{0} 1.01±0$$1\pm 0.002\times 10^{0}$ $1.02\pm 0.002\times 10^{0}$ $1.02\pm 0.002\times 10^{0}$ $1.03\pm 0.002\times 10^{0}$ $1.03\pm 0.002\times 10^{0}$
$(69.41\times)$ $(46.30\times)$ $(34.73\times)$ $(23.29\times)$ $(17.43\times)$ $(13.32\times)$
Random $2.74\pm 0.45\times 10^{-1}$ $5.59\pm 0.80\times 10^{-2}$ $1.33\pm 0.03\times 10^{-2}$ $9.06\pm 0.24\times 10^{-3}$ $6.87\pm 0.13\times 10^{-3}$ $5.66\pm 0.11\times 10^{-3}$
$(5.00\times)$ $(3.34\times)$ $(2.50\times)$ $(1.67\times)$ $(1.25\times)$ $(1.00\times)$
Active Learning $\textbf{9.32}\pm 6.96\times 10^{-2}$ $2.16\pm 0.66\times 10^{-2}$ $1.27\pm 0.03\times 10^{-2}$ $\textbf{7.86}\pm 0.14\times 10^{-3}$ $\textbf{6.24}\pm 0.21\times 10^{-3}$ $5.66\pm 0.11\times 10^{-3}$
$(5.00\times)$ $(3.33\times)$ $(2.50\times)$ $(1.67\times)$ $(1.25\times)$ $(1.00\times)$
[…]